Find the equation of the line that goes through the given point and has the given slope.
Write thea answer in slope-intercept form.
(2,5),4

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48565)

Answered
Cheers,
Stan H.

Answer by checkley71(8403)

Y=mX+b Y=5, X=2 & m=4 THUS
5=4*2+b
5=8+b
b=5-8
b=-3
THUS THE EQUATION IS
Y=4x-3